calculate the vapour pressure of solution at 25°C containing 26.9g of urea CO(NH2)2 in 712g H2O.( Vapour pressure of water at 25°C is 23.8mmHg)

1
Expert's answer
2022-05-31T11:26:04-0400

Solution:

The vapour pressure of pure water at 25°C is 23.8 mmHg


The molar mass of water (H2O) is 18.0153 g/mol

The molar mass of urea (CO(NH2)2) is 60.06 g/mol


Determine moles of water and urea:

Moles of water = (712 g water) × (1 mol water / 18.0153 g water) = 39.522 mol water

Moles of urea = (26.9 g urea) × (1 mol urea / 60.06 g urea) = 0.448 mol urea


Determine the mole fraction of the solvent (H2O):

χsolvent = (Moles of water) / (Moles of water + Moles of urea)

χsolvent = (39.522 mol) / (39.522 mol + 0.448 mol) = 0.9888

χsolvent = 0.9888


Using Raoult's Law, determine the vapour pressure of solution:

Psolution = χsolvent × P°solvent

Psolution = (0.9888) × (23.8 mmHg) = 23.53 mmHg

Psolution = 23.53 mmHg


Answer: The vapour pressure of solution is 23.53 mmHg